# Language Models Fail to Introspect About Their Knowledge of Language

### v2

@@ Acknowledgments
We thank Harvey Lederman and members of the UT Computational Linguistics Research Group for helpful comments and discussion. K.M. was supported by NSF grant 2139005. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
@@ Appendix A Format of metalinguistic prompts
As mentioned in § 3.1, we used both original and answer-reversed prompts ( and ) to eliminate bias caused by the answer position in our two experiments. The original prompts () for our experiments are presented in Table 2 (for acceptability judgments) and Table 3 (for word prediction). When provided with the original prompts(), the models choose the grammatical sentence (for sentence-level forced choice), the original continuation (for word-level forced choice) or ‘the sentence/word is good’ by responding with answer ‘1’. In reversed prompts, we change the positions of options, so the model should respond with answer ‘2’ to choose the answers (, or ‘Yes’) mentioned above. For instance, if of a minimal pair in Exp. 1 is:
@@
For instruction-tuned models, we use the default chat templates with the apply_chat_template function in Huggingface’s Transformers library (Wolf et al., 2019). We use the minicons (Misra, 2022) library to get model scores for calculating and . There is still debate over how models should be prompted to tackle multiple-choice questions in evaluation (Wang et al., 2024; Balepur et al., 2025). However, we believe our method is valid, as we used a number of different prompts and the models behave as expected when prompted, as shown in § 3.2.1 and § E.1.
@@ Appendix B Consistency between Direct and Meta evaluation methods
Fig. 5 shows the consistency (Cohen’s ) between Direct and Meta evaluation methods, for both Experiments 1 and 2. We use Cohen’s here to avoid overestimation of consistency, as both methods align well with some ground truth. As indicated by the low values, when we are getting forced choice answers (between two sentences or two words) from Direct and Meta methods, the consistency is low, even though the accuracy is high.
@@
In Exp. 1, larger models have better performance with the Meta method while all the models have similar accuracy with the Direct method (§ 3.2.1). Beyond this, we also found that scores from larger models tend to have higher agreement with scores in general, as shown by the brighter cells in the upper cells in Fig. 6, compared to the bottom ones.
@@
Similar to the results on the filtered dataset in Exp. 1(§ 3.2.2), we found no evidence of introspection when comparing Pearson values on unfiltered dataset, as shown in Fig. 7. Since our analysis is conducted on an unfiltered dataset, the differences between different ModelSimilarity appear to be smaller, and self does not have an advantage. We ran the same regression as in Exp. 1 that predicts from and ModelSimilarity and got similar results: there was a significant effect of (, ), and higher values of same family and other were observed (both ).
@@ E.1 Model preference on answers
Fig. 8 shows the proportion of times the model chooses the original last word rather than the randomly picked alternative word. For the wikipedia and news datasets, the models overwhelmingly prefer the original word (near 100% across all models), as it is both semantically and syntactically plausible. In contrast, for the nonsense and randomseq datasets where both continuations are randomly chosen and the sentences are meaningless, the model does not distinguish between the two options when prompted. We attribute the model’s preference for the original word in the nonsense dataset to its higher syntactic plausibility: both continuations ( and are meaningless, but follows some syntactic rules as its part of speech is the same as the final word in a grammatical sentence. For the randomseq dataset, although we controlled lemma frequency when selecting the alternative word, a significant difference in word-form frequency () was observed between the original word and alternative word (average log word form frequency=4.71 for and 5.31 for ). We believe this difference underlies the model’s slight preference for alternative words in the probability measurement.
@@ E.2 Effect of model size
Similar to Exp. 1, we found that from larger models from larger models tend to more correlated with from all the models across all the datasets, as shown by brighter upper cells in Fig. 9. We ran the same regression predicting mean Pearson from the log Direct model size, the Meta model size and their interaction. We found that the positive effect of Meta model size is significant (<0.001) across all datasets, the effect of interaction is insignificant and there is a negative effect of the Direct model size on wikipedia and news datasets.
@@ Appendix F Analysis on seed variants of the OLMo models
Fig. 10 shows a specific breakout for studying the OLMo models, the only models for which we have models that are identical besides their random seeds. We examined the within- and cross-model consistency of the 7B and 13B OLMo models and their variants. The three *B - seed* models for each parameter size are trained on the same data but with a different data order (different random seeds used), and the models “OLMo 7B” and “OLMo 13B” are final models with averaged weights. In these models trained on same data, there is still no sign of > . This pattern serves as strong support for our conclusion, that although is a function of ModelSimilarity, this does not indicate the introspection on knowledge of language. We ran regression models predicting with ‘whether the pair is self’ to examine the effect. Across all the datasets (filtered and unfiltered in Exp. 1, as well as four datasets in Exp. 2), no significant effect of self was observed (all ).
@@ Appendix G Regression models
For predicting model correctness from ModelSize and method (§ 3.2.1), we ran a logistic regression:
@@
In addition to using heatmaps to visualize the (Pearson ) between each pair of models (LABEL:fig:expr1-heatmap,Fig. 6, Fig. 9), we ran the following regression predicting with model size:
@@
Where mean_prompt_corr stands for Pearson (averaged across prompts), prompt_size stands for the size of the model and prob_size stands for the size of the model. Table 6 presents the regression models run on the results obtained from different datasets in this experiment.
@@
In addition to visualizing the relationship between (Pearson ) and ModelSimilarity (both categorical and empirical) in Fig. 3, Fig. 4 and Fig. 7, we ran the following regression predicting with both and categorical ModelSimilarity:
@@
Here, we use self as baseline and see the effects of other ModelSimilarity (seed variant, base/instruct, same family, other). Table 7 presents the regression models run on the results obtained from different datasets in this experiment.
@@
Table 7a shows on t the results we present in the main text, whereas Table 7b shows results where we focus in on just comparisons among “big models”, defined as those of 70B parameters or greater. As in the main text analysis, we would take a significant negative value for one of the ModelSimilarity coefficients to suggest an effect of introspection, but we do not find that. So, even among big models, we do not see evidence of introspection.
@@
As shown in Table 8, we do not observe any evidence of introspection among the OLMo models, as there is no significant effect of self across all datasets.

### v3

@@ Acknowledgments
We thank Harvey Lederman and members of the UT Computational Linguistics Research Group for helpful comments and discussion. K.M. was supported by NSF CAREER grant 2339729 from the Director of STEM Education (EDU) Division of Research on Learning in Formal and Informal (DRL) and OpenPhilanthropy funding to UT Austin. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
@@ Appendix A Format of metalinguistic prompts
As mentioned in §˜3.1, we used both original and answer-reversed prompts ( and ) to eliminate bias caused by the answer position in our two experiments. The original prompts () for our experiments are presented in Table˜2 (for acceptability judgments) and Table˜3 (for word prediction). When provided with the original prompts(), the models choose the grammatical sentence (for sentence-level forced choice), the original continuation (for word-level forced choice) or ‘the sentence/word is good’ by responding with answer ‘1’. In reversed prompts, we change the positions of options, so the model should respond with answer ‘2’ to choose the answers (, or ‘Yes’) mentioned above. For instance, if of a minimal pair in Exp. 1 is:
@@
For instruction-tuned models, we use the default chat templates with the apply_chat_template function in Huggingface’s Transformers library (Wolf et al., 2020). We use the minicons (Misra, 2022) library to get model scores for calculating and . There is still debate over how models should be prompted to tackle multiple-choice questions in evaluation (Wang et al., 2024; Balepur et al., 2025). However, we believe our method is valid, as we used a number of different prompts and the models behave as expected when prompted, as shown in §˜3.2.1 and §˜E.1.
@@ Appendix B Consistency between Direct and Meta evaluation methods
Fig.˜5 shows the consistency (Cohen’s ) between Direct and Meta evaluation methods, for both Experiments 1 and 2. We use Cohen’s here to avoid overestimation of consistency, as both methods align well with some ground truth. As indicated by the low values, when we are getting forced choice answers (between two sentences or two words) from Direct and Meta methods, the consistency is low, even though the accuracy is high.
@@
In Exp. 1, larger models have better performance with the Meta method while all the models have similar accuracy with the Direct method (§˜3.2.1). Beyond this, we also found that scores from larger models tend to have higher agreement with scores in general, as shown by the brighter cells in the upper cells in Fig.˜6, compared to the bottom ones.
@@
Similar to the results on the filtered dataset in Exp. 1(§˜3.2.2), we found no evidence of introspection when comparing Pearson values on unfiltered dataset, as shown in Fig.˜7. Since our analysis is conducted on an unfiltered dataset, the differences between different ModelSimilarity appear to be smaller, and self does not have an advantage. We ran the same regression as in Exp. 1 that predicts from and ModelSimilarity and got similar results: there was a significant effect of (, ), and higher values of same family and other were observed (both ).
@@ E.1 Model preference on answers
Fig.˜8 shows the proportion of times the model chooses the original last word rather than the randomly picked alternative word. For the wikipedia and news datasets, the models overwhelmingly prefer the original word (near 100% across all models), as it is both semantically and syntactically plausible. In contrast, for the nonsense and randomseq datasets where both continuations are randomly chosen and the sentences are meaningless, the model does not distinguish between the two options when prompted. We attribute the model’s preference for the original word in the nonsense dataset to its higher syntactic plausibility: both continuations ( and are meaningless, but follows some syntactic rules as its part of speech is the same as the final word in a grammatical sentence. For the randomseq dataset, although we controlled lemma frequency when selecting the alternative word, a significant difference in word-form frequency () was observed between the original word and alternative word (average log word form frequency=4.71 for and 5.31 for ). We believe this difference underlies the model’s slight preference for alternative words in the probability measurement.
@@ E.2 Effect of model size
Similar to Exp. 1, we found that from larger models from larger models tend to more correlated with from all the models across all the datasets, as shown by brighter upper cells in Fig.˜9. We ran the same regression predicting mean Pearson from the log Direct model size, the Meta model size and their interaction. We found that the positive effect of Meta model size is significant (<0.001) across all datasets, the effect of interaction is insignificant and there is a negative effect of the Direct model size on wikipedia and news datasets.
@@ Appendix F Analysis on seed variants of the OLMo models
Fig.˜10 shows a specific breakout for studying the OLMo models, the only models for which we have models that are identical besides their random seeds. We examined the within- and cross-model consistency of the 7B and 13B OLMo models and their variants. The three *B - seed* models for each parameter size are trained on the same data but with a different data order (different random seeds used), and the models “OLMo 7B” and “OLMo 13B” are final models with averaged weights. In these models trained on same data, there is still no sign of > . This pattern serves as strong support for our conclusion, that although is a function of ModelSimilarity, this does not indicate the introspection on knowledge of language. We ran regression models predicting with ‘whether the pair is self’ to examine the effect. Across all the datasets (filtered and unfiltered in Exp. 1, as well as four datasets in Exp. 2), no significant effect of self was observed (all ).
@@ Appendix G Regression models
For predicting model correctness from ModelSize and method (§˜3.2.1), we ran a logistic regression:
@@
In addition to using heatmaps to visualize the (Pearson ) between each pair of models (LABEL:fig:expr1-heatmap,Fig.˜6, Fig.˜9), we ran the following regression predicting with model size:
@@
Where mean_prompt_corr stands for Pearson (averaged across prompts), prompt_size stands for the size of the model and prob_size stands for the size of the model. Table˜6 presents the regression models run on the results obtained from different datasets in this experiment.
@@
In addition to visualizing the relationship between (Pearson ) and ModelSimilarity (both categorical and empirical) in Fig.˜3, Fig.˜4 and Fig.˜7, we ran the following regression predicting with both and categorical ModelSimilarity:
@@
Here, we use self as baseline and see the effects of other ModelSimilarity (seed variant, base/instruct, same family, other). Table˜7 presents the regression models run on the results obtained from different datasets in this experiment.
@@
Table˜7a shows on t the results we present in the main text, whereas Table˜7b shows results where we focus in on just comparisons among “big models”, defined as those of 70B parameters or greater. As in the main text analysis, we would take a significant negative value for one of the ModelSimilarity coefficients to suggest an effect of introspection, but we do not find that. So, even among big models, we do not see evidence of introspection.
@@
As shown in Table˜8, we do not observe any evidence of introspection among the OLMo models, as there is no significant effect of self across all datasets.